OpenStudy (anonymous):

How would I factor: 3x^2 - x - 2 = 0

OpenStudy (callisto):

3x^2 - x - 2 = 0 3x^2 - 3x + 2x -2 =0 3x(x-1) +2(x-1) =0 (3x+2)(x-1)=0

OpenStudy (anonymous):

what happens in that second step?

OpenStudy (callisto):

Hmm.. just a little trick... since -x = -3x+2x This helps factoring by grouping

OpenStudy (anonymous):

ahh i think i see

OpenStudy (callisto):

Actually, my way of doing factorization is this:  But it seems more difficult to understand :|